# arXiv:2602.23780v1[eess.SP] 27 Feb 2026

### 27 Feb 2026 v1

@@ Abstract
Convolution serves as a powerful operation for the regularization of functions. While polynomials inherently possess smoothness, it is particularly interesting to investigate their behavior under convolution. This interest stems from the fact that numerous engineering and physical phenomena can be modeled through such operations, including weighted averages, blurring effects and convolutional integral equations. In this work, we show that under certain mild conditions, the convolution with any even Schwartz function acts as an automorphism on the vector space of finite-order polynomials. We derive explicit equations for the inverse operation of this convolution, which are numerically simple to implement. In addition, we extend the deconvolution with (not necessarily even) Schwartz functions to a broader class of functions, including $L^{1}(\mathbb{R})$, $L^{2}(\mathbb{R})$, the Schwartz space and tempered distributions. Specifically, we establish a explicit rigorous formula for the deconvolution of a function or distribution that has been convolved with a Schwartz function, being a particular example the Weierstrass Transform. For the latter, we show that any Schwartz function and tempered distribution that has been transformed, can be recovered, in their respective topologies, by the limit of a sequence of linear combination of recursive convolutions. This provides a new formula for the inverse of the Weierstrass Transform that can be numerically implemented.
@@
Deconvolution, Weierstrass Transform, Polynomials, Schwartz Functions
pacs:
MSC Classification]44A35, 45E10, 42A85, 46F12, 44A15
@@ 1 Introduction
where $x\mapsto\varphi_{\varepsilon}(x):=\frac{1}{\varepsilon\sqrt{4\pi}}e^{-\frac{x^{2}}{4\varepsilon^{2}}}$ is the so called Gaussian Kernel. The Gaussian Kernel is an even Schwartz function that satisfies the following properties. It has unit area, i.e., $\int_{\mathbb{R}}\varphi_{\varepsilon}=1$, it converges to the Dirac delta distribution as $\varepsilon\to 0^{+}$, and for any $f\in L^{p}_{loc}(\mathbb{R})$, $F_{\varepsilon}\in C^{\infty}(\mathbb{R})$ and $F_{\varepsilon}\to f$ as $\varepsilon\to 0^{+}$ in several types of convergence. In particular it $f\in L^{p}_{loc}(\mathbb{R})$ then $F_{\varepsilon}\to f$ as $\varepsilon\to 0^{+}$ pointwise and in $L^{p}$. It has been studied extensively in [widder1951necessary], where its inversion is presented as a differential operator, but also in [bilodeau1962weierstrass] where it is shown the relationship between the transform and the Hermite polynomials.
Observing the inversion formula in the aforementioned references, and how the Weierstrass transform acts on polynomials, it is natural to ask how the polynomials behave under a more general operation, and if there is an inverse for this operation which is simple to implement and compute That is, instead of focusing on the convolution with the Gaussian Kernel, we focus on the convolution with Schwartz convolution kernels. These questions, apart from its theoretical interest have interesting applications, since convolution with a function of the above characteristics can be thought as a weighted average operation, or equivalently, as a low pass filter, and so the process of deconvolution, i.e., the process of restoring an original signal from a filtered one, is of great application and theoretical interest. The theoretical importance of deconvolution is well-established. Classical treatments such as [berenstein1983some] and [hirschman2012convolution] provide systematic inversion formulas for several convolution transforms, [hohlfeld1993solution] provides explicit formulas for the solution of convolution integral equations by means of differentiation and [prost2003deconvolution] presents a formula for inverting the convolution when the convolution kernel has a Fourier transform with compact support. Moreover, in [baeumer2003inversion] is presented an inverse convolution approach by the limit of a finite sequence of linear combinations of exponential functions. More recently, in [kaiser2025deconvolutionarbitrarydistributionfunctions] deconvolution formulas for arbitrary distribution functions and densities using characteristic functions and geometric series methods are developed. Beyond theory, deconvolution has several practical applications and implementations. Numerical approaches are deeply treated in [Hansen2002-li], while [madden1996comparison] provides a comparative analysis of different deconvolution techniques. In [ulmer2010inverse] the deconvolution with Gaussian Kernels is treated and applied to radiotherapy problems, and in [fanton2021convolution] it is discussed several physical and engineering applications where (de)convolution is necessary. If the convolution kernel is unknown, blind deconvolution techniques become useful, as presented in [campisi2017blind] and [zhang2023fast]. Convolutions are also applied in more mathematical abstract situations, as in [cech2025optimalitymollifiers]. More recently, [Hohage_Marechal_Simar_Vanhems_2024] provides a novel framework for deconvolution techniques.
@@
In this work we contribute both theoretically and practically to the deconvolution field with Theorem 2 and, particularly, with Theorem 3. In the former we present that the convolution with an even Schwartz function is an invertible operator on the space of finite order polynomials. In the latter we show that the inverse of the convolution with a Schwartz function can be explicitly computed as a limit of a sequence of linear combinations of recursive convolutions. More specifically, in Section 2 we show that, if the convolution kernel is an even Schwartz function, the convolution operator is an automorphism of polynomial vector spaces of finite degree. Moreover, we explicitly present the formula (Equations (7) and (8)) of the inverse of the convolution, and validate the results numerically. The inverse, curiously enough, consists on a linear combination of iterative convolutions of the convolved polynomial (see Theorem 2), from which it follows that any polynomial can be written as a linear combination of its iterative convolutions; that is, any polynomial is, in essence, a linear combination of its weighted averages, see Remark 4. We end the section discussing the main result (Theorem 3), which we consider the main contribution to the field of deconvolution, either theoretically and numerically speaking. It happens that, by passing to the limit in Theorem 2, we obtain Theorem 3, that is, we can recover as a limiting sequence of linear combinations of recursive convolutions the original function which has been convolved, obtaining, particularly, a new formula for the inverse of the Weierstrass Transform. For this special case of the Weierstrass transform, we also prove that any Schwartz function or tempered distribution can be recovered by passing to the limit in Equation (9), in their respective topologies. Clearly, several assumptions must be made clear for the original function and the convolution kernel, which do not contradict the traditional assumptions for the inverse of the convolution to exist in the Fourier transformed space. In Section 3 we present some applications of these results, validating our approach and providing some insight for new signal processing methods. We end this work with some conclusions in Section 4.
@@ 1.1 Notation
For a set $A$ we denote its closure as $\overline{A}$. The support of a real-valued function is defined as $\operatorname{supp}f:=\overline{\{x\in\operatorname{dom}f\mid f(x)\neq 0\}}$, where $\operatorname{dom}f$ is the domain of $f$. Given $x,y\in\mathbb{R}$ we define $xA+y=\{xa+y\mid x\in A\}$. If $n>2$ a natural number, by $\sum_{j\in 2\mathbb{N}_{0}}^{n}j$ we mean the sum of even nonnegative integers up to $n$ if $n$ is even, or up to n1n-1 if $n$ is odd. The sum $\sum_{j\in 2\mathbb{N}}^{n}j$ is then understood as before but for elements in $2\mathbb{N}$. Given two functions $f,g:\mathbb{R}^{d}\to\mathbb{R}$ with $d\in\mathbb{N}$ we denote, when it exists, its convolution as fgf*g. The space of Schwartz functions is denoted as $\mathscr{S}(\mathbb{R}^{d})$, while $\mathscr{S}_{M}(\mathbb{R}^{d})=\{f\in\mathscr{S}(\mathbb{R}^{d})\mid\int_{\mathbb{R}^{d}}f=1\}$, and $\mathscr{S}^{\prime}(\mathbb{R}^{d})$ denotes the space of tempered distributions. For $f\in\mathscr{S}^{\prime}(\mathbb{R}^{d})$ and $\psi\in\mathscr{S}(\mathbb{R}^{d})$ we denote the image of $\psi$ under $f$ either by $f(\psi)$ or $\langle f,\psi\rangle$. The space of $p\in[1,\infty]$ integrable functions as $L^{p}(\mathbb{R}^{d})$, and the space of locally integrable functions as $L^{p}_{loc}(\mathbb{R}^{d})$, i.e., $p$-integrable on each compact subset of $\mathbb{R}^{d}$. For a function $\varphi:\mathbb{R}^{d}\to\mathbb{R}$ we will denote by $\varphi_{\varepsilon}:=\frac{1}{\varepsilon^{d}}\varphi\circ\frac{\operatorname{id}}{\varepsilon}$ for any $\varepsilon>0$. We define the Fourier transform of a function $f:\mathbb{R}\to\mathbb{R}$, when it exists, by $\mathcal{F}(f)(\xi)=\hat{f}(\xi):=\int_{\mathbb{R}}f(x)e^{-2\pi i\xi x}dx$ for all $\xi\in\mathbb{R}$, and its inverse Fourier transform by $\mathcal{F}^{-1}(f)$. By LDCT we refer to the Lebesgue’s Dominated Convergence Theorem and by FFT we mean the Fast Fourier Transform.
@@
  • If $\varphi$ is even, the convolution of $p$ with $\varphi$ results in a polynomial of the same degree (Theorem 1). In particular, if the original degree of the polynomial is $n\geq 2$, the convolution is the sum of the original polynomial with another polynomial of degree n2n-2, one or zero, depending on the parity of $n$ (Corollaries 1 and 2). In particular, the convolution with an even Schwartz function of a polynomial is another polynomial with the same asymptotic growth, and in addition the first and second leading coefficients are shared between $p$ and its convolution.
  • If $\varphi$ is even, for any $\varepsilon>0$, the convolution operator $p\mapsto p*\varphi_{\varepsilon}$ is an automorphism between the vector space of polynomials of the same degree, for any value of $\varepsilon>0$ (Theorem 2). This implies that the convolution with an even Schwartz kernel in the vector space of polynomials is invertible. The inverse can be computed as in Equations (7) and (8), and we provide an insight of its numerical simplicity in Corollary 3.
  • We prove in Theorem 3, that these results can be generalized by passing to the limit in Equation (7) of Theorem 2, for a certain class of functions, and without the need for the Schwartz convolution kernel to be even. This results are remarkable, since by an iterative process the original function can be restored via convolutions.
@@ 2.1 Polynomials in the real line
$\int_{\mathbb{R}}(x-y)^{m}\varphi_{\varepsilon}(y)dy$ $=\int_{\mathbb{R}}\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}x^{m-k}y^{k}\varphi_{\varepsilon}(y)dy$
$=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}x^{m-k}\int_{\mathbb{R}}y^{k}\varphi_{\varepsilon}(y)dy$
@@
where the last equality comes from Definition 1 and the fact that $(-1)^{k}=1$ for even $k$. Now note that since the convolution is a bilinear operation
@@
and since the sums are finite, we can use Fubini’s Theorem to end the proof of Equation (3). Note that the right hand side of Equation (3) is again a polynomial of at most degree $n$, proving the theorem. ∎
It is clear and immediate that if $\varphi$ is not even we still get a polynomial, the fact that we are restricting to even convolution kernels will become apparent in Theorem 2.
We now can improve upon the results of Theorem 1 using the following corollary.
Corollary 1.
@@
That is, the convolution of a polynomial results in another polynomial with the same asymptotic growth, in fact the first two leading coefficients of $P_{\varepsilon}$ is the same as $p$.
Proof.
We proceed for the even case by induction over $n\geq 2$. Suppose first that $n=2$, then it is straightforward to see using Theorem 1 that
@@
and so Equation (4) is valid for $n=2$. Let’s suppose it is true for $n\in\mathbb{N}$ and proceed to prove that it is also true for $n+1$. Let $q\in\mathscr{P}_{n+1}(\mathbb{R})$, and denote by $r\in\mathscr{P}_{n}(\mathbb{R})$ the polynomial of degree $n$ obtained by making the leading coefficient of $q$ to be zero. Then by Theorem 1,
$Q_{\varepsilon}(x)=(q*\varphi_{\varepsilon})(x)=\sum_{j=0}^{n+1}\sum_{k\in 2\mathbb{N}_{0}}^{j}\binom{j}{k}a_{j}x^{j-k}c_{k}\varepsilon^{k}$
@@
$\sum_{k\in 2\mathbb{N}}^{n+1}\binom{n+1}{k}a_{n+1}x^{n+1-k}c_{k}\varepsilon^{k}=x$ $^{n-1}\binom{n+1}{2}c_{2}\varepsilon^{2}a_{n+1}+x^{n-3}\binom{n+1}{4}c_{4}\varepsilon^{4}a_{n+1}$
$+\dots+x\binom{n+1}{n}c_{n}\varepsilon^{n}a_{n+1},$
@@
$\sum_{j=0}^{n-2}x^{j}\left(\sum_{k\in 2\mathbb{N}}^{n-j}\binom{k+j}{k}c_{k}\varepsilon^{k}a_{k+j}\right)+\sum_{k\in 2\mathbb{N}}^{n+1}\binom{n+1}{k}a_{j}x^{n+1-k}c_{k}\varepsilon^{k}$
$=\sum_{j=0}^{n-1}x^{j}\left(\sum_{k\in 2\mathbb{N}}^{n+1-j}\binom{k+j}{k}c_{k}\varepsilon^{k}a_{k+j}\right),$
completing the proof for even $n$. If $n$ is odd the proof is similar. ∎
Remark 2.
@@
where $p_{j}$ is defined recursively as follows. If $j=0$ then $p_{0}:=p$. If $j=1$ then $p_{j}:=\tilde{p}$ as in Remark 2, if $j=2$ then $p_{j}:=\tilde{\tilde{p}}$, which comes from $T(\tilde{p})=\tilde{p}+\tilde{\tilde{p}}$, as in Remark 2, and so on. Moreover the following statements are also true.
@@
where $p_{1}\in\mathscr{P}_{n-2}(\mathbb{R})$ by Corollary 1. Note that $T$ is linear, which implies
@@
So the statement is proven for $k=2$, because $p_{1}\in\mathscr{P}_{n-2}(\mathbb{R})$ by Corollary 1 and Remark 2, which by the same conclusions imply that $p_{2}\in\mathscr{P}_{\max\{n-4,0\}}(\mathbb{R})$.
@@
$T_{\varepsilon}^{k+1}(p)$ $=\sum_{j=1}^{k}\left(\binom{k}{j}+\binom{k}{j-1}\right)p_{j}+p_{0}+p_{k+1}$
$=\sum_{j=1}^{k}\binom{k+1}{j}p_{j}+p_{0}+p_{k+1}=\sum_{j=0}^{k+1}\binom{k+1}{j}p_{j},$
where we used the fact that $\binom{k}{j}+\binom{k}{j-1}=\binom{k+1}{j}$. Moreover, $p_{k}\in\mathscr{P}_{n-2k}(\mathbb{R})$, which implies by Corollary 1 and Remark 2 that $p_{k+1}\in\mathscr{P}_{n-2(k+1)}(\mathbb{R})$. This proves Equation (5). Note that the formula is always valid, independently if $n$ is even or odd, so what we need now is to prove it for when $2k>n$.
@@
It will be useful to compute the polynomials $p_{j}$ in Corollary 2 as linear combinations of repeated convolutions of $p_{0}=p$, i.e., of the original polynomial. This will allow to obtain a formula for the inversion of the convolution on the space $\mathscr{P}_{n}(\mathbb{R})$. We prove that this is possible in the next lemma.
Lemma 1.
Let $p\in\mathscr{P}_{n}(\mathbb{R})$ with $n\geq 2$, $\varphi\in\mathscr{S}_{M}(\mathbb{R})$ be even and $\varepsilon>0$. Defining $T_{\varepsilon}:\mathscr{P}_{n}(\mathbb{R})\to\mathscr{P}_{n}(\mathbb{R})$ as in Corollary 2 as well as $p_{j}$ for $j\in\mathbb{N}_{0}$, then
@@
Proof.
We proceed by induction over $j$. By definition it is true for $j=0$, and for $j=1$ we know by Theorem 1 and Corollary 2 that $T_{\varepsilon}(p_{0})=p_{0}+p_{1}$ and therefore $p_{1}=T_{\varepsilon}(p_{0})-p_{0}$ for which Equation (6) clearly holds. Suppose now that it is true for $j\in\mathbb{N}$. Then $T_{\varepsilon}(p_{j})=p_{j}+p_{j+1}$ and so

### 17 Apr 2026 v2

@@ Abstract
In this work, we construct an explicit, theoretically rigorous deconvolution method that relies entirely on iterative forward convolutions, thus can be numerically implemented. We first prove that convolution with an even Schwartz kernel acts as an automorphism on the vector space of finite-degree polynomials. Exploiting the parity of the kernel, we derive an exact algebraic inverse for this space, expressed uniquely as a finite linear combination of repeated convolutions. The core contribution of this work extends this algebraic inversion to infinite-dimensional function spaces, including $L^{1}(\mathbb{R})$, $L^{2}(\mathbb{R})$, the Schwartz space $\mathscr{S}(\mathbb{R})$, and the space of tempered distributions $\mathscr{S}^{\prime}(\mathbb{R})$. By passing the finite-sum polynomial inversion formula to the limit, we demonstrate that an arbitrary function or distribution convolved with a Schwartz kernel can be exactly recovered in its respective topology. The resulting inverse is an explicitly computable limit of a sequence of linear combinations of recursive convolutions. As a primary application, this limit provides a fundamentally new, iterative numerical formula for the inverse of the Weierstrass Transform. By bypassing traditional numerically ill-posed inversion techniques, our method offers a mathematically exact and numerically robust algorithm for computational signal recovery.
@@
[…]
{highlights}
Convolution with an even Schwartz kernel is a polynomial automorphism.
Exact polynomial deconvolution via finite iterative forward convolutions.
Recursive limits extend exact inversion to $L^{1}$, $L^{2}$, and distributions.
Bypasses ill-posed techniques, recovering functions in their topologies.
Yields a robust iterative numerical inverse for the Weierstrass transform.
@@ 1 Introduction
where $x\mapsto\varphi_{\varepsilon}(x):=\frac{1}{\varepsilon\sqrt{4\pi}}e^{-\frac{x^{2}}{4\varepsilon^{2}}}$ is the so called Gaussian Kernel. The Gaussian Kernel is an even Schwartz function that satisfies several properties. It has unit area, i.e., $\int_{\mathbb{R}}\varphi_{\varepsilon}=1$, it converges to the Dirac delta distribution as $\varepsilon\to 0^{+}$, and for any $f\in L^{p}_{loc}(\mathbb{R})$, $F_{\varepsilon}\in C^{\infty}(\mathbb{R})$ and $F_{\varepsilon}\to f$ as $\varepsilon\to 0^{+}$ in several types of convergence. In particular if $f\in L^{p}_{loc}(\mathbb{R})$ then $F_{\varepsilon}\to f$ as $\varepsilon\to 0^{+}$ pointwise and in $L^{p}$. The theoretical foundation of this transform has been studied extensively in widder1951necessary, where its inversion is presented as a differential operator, and in bilodeau1962weierstrass, which explores the relationship between the transform and Hermite polynomials.
While classical inversion formulas provide elegant theoretical solutions, they often rely on unbounded differential operators or Fourier domain divisions that can be highly susceptible to numerical instability and noise. Observing the action of the Weierstrass transform on simple spaces like polynomials, it is natural to ask how polynomials behave under a more general convolution kernel, and crucially, if there is an inverse for this operation which is both mathematically rigorous and simple to compute numerically which can be extended to a bigger space of functions. That is, instead of focusing solely on the Gaussian Kernel, we generalize our study to convolution with arbitrary Schwartz kernels. These questions possess deep computational applications, since convolution with such a function models weighted averages, low-pass filters, and various physical phenomena. Consequently, the process of deconvolution — the restoration of an original signal from a filtered one — is an important problem in computational and applied mathematics.
The theoretical and computational importance of deconvolution is well-established across applied mathematics. Classical treatments such as berenstein1983some and hirschman2012convolution provide systematic inversion formulas for various convolution transforms, while explicit solutions to convolution integral equations have historically relied on differentiation hohlfeld1993solution or Fourier domain restrictions prost2003deconvolution. To overcome the numerical vulnerabilities of these classical analytical methods, modern approaches often utilize limiting sequences baeumer2003inversion, geometric series methods for arbitrary distribution functions kaiser2025deconvolutionarbitrarydistributionfunctions, and regularized iterative filtering techniques LI2018425. Furthermore, in finite-dimensional discrete settings, structural algebraic frameworks — such as polynomial-based matrix approaches — have been successful in performing exact deconvolution DIAZTOCA201798. However, when transitioning to practical numerical implementations, carefully managing convolution matrices and boundary conditions remains a significant computational challenge ZHOU201414; Hansen2002-li. This underlying complexity has prompted comparative numerical analyses madden1996comparison and continues to drive modern blind deconvolution techniques campisi2017blind; zhang2023fast; XU2025116511; DYKES2021113067 and broader algorithmic frameworks Hohage_Marechal_Simar_Vanhems_2024 across diverse physical and medical applications ulmer2010inverse; fanton2021convolution; HOWARD201443.
@@
In this work, we contribute to this field by developing a novel, computationally viable framework for exact deconvolution. In Section 2, we prove that convolution with an even Schwartz function is an invertible operator (an automorphism) on the space of finite-order polynomials (Theorem 2.12). Then, instead of taking an algebraic perspective, we also show that the exact inverse of this operator can be expressed purely as a finite linear combination of iterative convolutions (Equations (7) and (8)). The main contribution of this paper, both theoretically and computationally, is the extension of this recursive inversion framework to broader function spaces (Theorem 2.18). By passing the polynomial inversion formula to the limit, we explicitly recover the original function or tempered distribution as a limiting sequence of linear combinations of recursive convolutions. This yields a novel, iterative numerical formula for the inverse of the Weierstrass Transform. We rigorously prove convergence in the respective topologies ($L^{1}$, $L^{2}$, the Schwartz space, and tempered distributions), ensuring that the method is theoretically correct while remaining numerically implementable. Finally, in Section 3, we present applications of these results, validating our approach and providing numerical insight into how this iterative convolution framework translates into robust signal processing algorithms. We end this work with conclusions in Section 4.
@@ 1.1 Notation
For a set $A$ we denote its closure as $\overline{A}$. The support of a real-valued function is defined as
$\operatorname{supp}f:=\overline{\{x\in\operatorname{dom}f\mid f(x)\neq 0\}},$
where $\operatorname{dom}f$ is the domain of $f$. Given $x,y\in\mathbb{R}$ we define $xA+y=\{xa+y\mid x\in A\}$. If $n>2$ a natural number, by $\sum_{j\in 2\mathbb{N}_{0}}^{n}j$ we mean the sum of even nonnegative integers up to $n$ if $n$ is even, or up to n1n-1 if $n$ is odd. The sum $\sum_{j\in 2\mathbb{N}}^{n}j$ is then understood as before but for elements in $2\mathbb{N}$. Given two functions $f,g:\mathbb{R}^{d}\to\mathbb{R}$ with $d\in\mathbb{N}$ we denote, when it exists, its convolution as fgf*g. The space of Schwartz functions is denoted as $\mathscr{S}(\mathbb{R}^{d})$, while $\mathscr{S}_{M}(\mathbb{R}^{d})=\{f\in\mathscr{S}(\mathbb{R}^{d})\mid\int_{\mathbb{R}^{d}}f=1\}$, and $\mathscr{S}^{\prime}(\mathbb{R}^{d})$ denotes the space of tempered distributions. For $f\in\mathscr{S}^{\prime}(\mathbb{R}^{d})$ and $\psi\in\mathscr{S}(\mathbb{R}^{d})$ we denote the image of $\psi$ under $f$ either by $f(\psi)$ or $\langle f,\psi\rangle$. The space of $p\in[1,\infty]$ integrable functions as $L^{p}(\mathbb{R}^{d})$, and the space of locally integrable functions as $L^{p}_{loc}(\mathbb{R}^{d})$, i.e., $p$-integrable on each compact subset of $\mathbb{R}^{d}$. For a function $\varphi:\mathbb{R}^{d}\to\mathbb{R}$ we will denote by $\varphi_{\varepsilon}:=\frac{1}{\varepsilon^{d}}\varphi\circ\frac{\operatorname{id}}{\varepsilon}$ for any $\varepsilon>0$. We define the Fourier transform of a function $f:\mathbb{R}\to\mathbb{R}$, when it exists, by $\mathcal{F}(f)(\xi)=\hat{f}(\xi):=\int_{\mathbb{R}}f(x)e^{-2\pi i\xi x}dx$ for all $\xi\in\mathbb{R}$, and its inverse Fourier transform by $\mathcal{F}^{-1}(f)$. By LDCT we refer to the Lebesgue’s Dominated Convergence Theorem and by FFT we mean the Fast Fourier Transform.
@@
  • If $\varphi$ is even, the convolution of $p$ with $\varphi$ results in a polynomial of the same degree (Theorem 2.3). In particular, if the original degree of the polynomial is $n\geq 2$, the convolution is the sum of the original polynomial with another polynomial of degree n2n-2, one or zero, depending on the parity of $n$ (Corollaries 2.5 and 2.8). In particular, the convolution with an even Schwartz function of a polynomial is another polynomial with the same asymptotic growth, and in addition the first and second leading coefficients are shared between $p$ and its convolution.
  • If $\varphi$ is even, for any $\varepsilon>0$, the convolution operator $p\mapsto p*\varphi_{\varepsilon}$ is an automorphism between the vector space of polynomials of the same degree, for any value of $\varepsilon>0$ (Theorem 2.12). This implies that the convolution with an even Schwartz kernel in the vector space of polynomials is invertible. The inverse can be computed as in Equations (7) and (8), and we provide an insight of its numerical simplicity in Corollary 2.14.
  • We prove in Theorem 2.18, that these results can be generalized by passing to the limit in Equation (7) of Theorem 2.12, for a certain class of functions, and without the need for the Schwartz convolution kernel to be even. This results are remarkable, since by an iterative process the original function can be restored via convolutions.
@@ 2.1 Polynomials in the real line
$\int_{\mathbb{R}}(x-y)^{m}\varphi_{\varepsilon}(y)dy$ $=\int_{\mathbb{R}}\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}x^{m-k}y^{k}\varphi_{\varepsilon}(y)dy\$$=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}x^{m-k}\int_{\mathbb{R}}y^{k}\varphi_{\varepsilon}(y)dy$
@@
where the last equality comes from Definition 2.1 and the fact that $(-1)^{k}=1$ for even $k$. Now note that since the convolution is a bilinear operation
@@
and since the sums are finite, we can use Fubini’s Theorem to end the proof of Equation (3). Note that the right hand side of Equation (3) is again a polynomial of at most degree $n$, proving the theorem.
It is clear and immediate that if $\varphi$ is not even we still get a polynomial, the fact that we are restricting to even convolution kernels will become apparent in Theorem 2.12.
We now can improve upon the results of Theorem 2.3 using the following corollary.
Corollary 2.5.
@@
That is, the convolution of a polynomial results in another polynomial with the same asymptotic growth, in fact the first two leading coefficients of $P_{\varepsilon}$ are the same as those of $p$.
Proof 2.6.
We proceed for the even case by induction over $n\geq 2$. Suppose first that $n=2$, then it is straightforward to see using Theorem 2.3 that
@@
and so Equation (4) is valid for $n=2$. Let’s suppose it is true for $n\in\mathbb{N}$ and proceed to prove that it is also true for $n+1$. Let $q\in\mathscr{P}_{n+1}(\mathbb{R})$, and denote by $r\in\mathscr{P}_{n}(\mathbb{R})$ the polynomial of degree $n$ obtained by making the leading coefficient of $q$ to be zero. Then by Theorem 2.3,
$Q_{\varepsilon}(x)$ $:=(q*\varphi_{\varepsilon})(x)=\sum_{j=0}^{n+1}\sum_{k\in 2\mathbb{N}_{0}}^{j}\binom{j}{k}a_{j}x^{j-k}c_{k}\varepsilon^{k}$
@@
$\sum_{k\in 2\mathbb{N}}^{n+1}\binom{n+1}{k}a_{n+1}x^{n+1-k}c_{k}\varepsilon^{k}=x$ $^{n-1}\binom{n+1}{2}c_{2}\varepsilon^{2}a_{n+1}+x^{n-3}\binom{n+1}{4}c_{4}\varepsilon^{4}a_{n+1}+\dots+x\binom{n+1}{n}c_{n}\varepsilon^{n}a_{n+1},$
@@
$\sum_{j=0}^{n-2}x^{j}\left(\sum_{k\in 2\mathbb{N}}^{n-j}\binom{k+j}{k}c_{k}\varepsilon^{k}a_{k+j}\right)+\sum_{k\in 2\mathbb{N}}^{n+1}\binom{n+1}{k}a_{j}x^{n+1-k}c_{k}\varepsilon^{k}=\sum_{j=0}^{n-1}x^{j}\left(\sum_{k\in 2\mathbb{N}}^{n+1-j}\binom{k+j}{k}c_{k}\varepsilon^{k}a_{k+j}\right),$
completing the proof for even $n$. If $n$ is odd the proof is similar.
Remark 2.7.
@@
where $p_{j}$ is defined recursively as follows. If $j=0$ then $p_{0}:=p$. If $j=1$ then $p_{j}:=\tilde{p}$ as in Remark 2.7, if $j=2$ then $p_{j}:=\tilde{\tilde{p}}$, which comes from $T(\tilde{p})=\tilde{p}+\tilde{\tilde{p}}$, as in Remark 2.7, and so on. Moreover the following statements are also true.
@@
where $p_{1}\in\mathscr{P}_{n-2}(\mathbb{R})$ by Corollary 2.5. Note that $T$ is linear, which implies
@@
So the statement is proven for $k=2$, because $p_{1}\in\mathscr{P}_{n-2}(\mathbb{R})$ by Corollary 2.5 and Remark 2.7, which by the same conclusions imply that $p_{2}\in\mathscr{P}_{\max\{n-4,0\}}(\mathbb{R})$.
@@
$T_{\varepsilon}^{k+1}(p)$ $=\sum_{j=1}^{k}\left(\binom{k}{j}+\binom{k}{j-1}\right)p_{j}+p_{0}+p_{k+1}=\sum_{j=1}^{k}\binom{k+1}{j}p_{j}+p_{0}+p_{k+1}=\sum_{j=0}^{k+1}\binom{k+1}{j}p_{j},$
where we used the fact that $\binom{k}{j}+\binom{k}{j-1}=\binom{k+1}{j}$. Moreover, $p_{k}\in\mathscr{P}_{n-2k}(\mathbb{R})$, which implies by Corollary 2.5 and Remark 2.7 that $p_{k+1}\in\mathscr{P}_{n-2(k+1)}(\mathbb{R})$. This proves Equation (5). Note that the formula is always valid, independently if $n$ is even or odd, so what we need now is to prove it for when $2k>n$.
@@
It will be useful to compute the polynomials $p_{j}$ in Corollary 2.8 as linear combinations of repeated convolutions of $p_{0}=p$, i.e., of the original polynomial. This will allow to obtain a formula for the inversion of the convolution on the space $\mathscr{P}_{n}(\mathbb{R})$. We prove that this is possible in the next lemma.
Lemma 2.10.
Let $p\in\mathscr{P}_{n}(\mathbb{R})$ with $n\geq 2$, $\varphi\in\mathscr{S}_{M}(\mathbb{R})$ be even and $\varepsilon>0$. Defining $T_{\varepsilon}:\mathscr{P}_{n}(\mathbb{R})\to\mathscr{P}_{n}(\mathbb{R})$ as in Corollary 2.8 as well as $p_{j}$ for $j\in\mathbb{N}_{0}$, then
@@
Proof 2.11.
We proceed by induction over $j$. By definition it is true for $j=0$, and for $j=1$ we know by Theorem 2.3 and Corollary 2.8 that $T_{\varepsilon}(p_{0})=p_{0}+p_{1}$ and therefore $p_{1}=T_{\varepsilon}(p_{0})-p_{0}$ for which Equation (6) clearly holds. Suppose now that it is true for $j\in\mathbb{N}$. Then $T_{\varepsilon}(p_{j})=p_{j}+p_{j+1}$ and so